Prove that a non-decreasing (resp. non-increasing) sequence which is not

bounded above (resp. bounded below) diverges to +∞ (resp. to −∞).


1
Expert's answer
2021-02-24T12:26:56-0500

Solution:

To prove: A non-increasing sequence which is not bounded below diverges to −infinity.

Proof: Let "\left\{a_{n}\right\}" be a monotonically non-increasing sequence.

"\Rightarrow \quad a_{n+1}<a_{n} \forall n \in {N}"

"\left\{a_{n}\right\}" is not bounded below.

"\Rightarrow \quad" For any "M \in {R}, \exists m \in {N}" such that "a_{k}<M \forall k>m" .

"\Rightarrow \quad" "\lim _{n \rightarrow \infty} a_{n}=-\infty"

"\Rightarrow \quad" A monotonically decreasing sequence that is not bounded below diverges to negative infinity.

Similarly, we can do other part.

To prove: A non-decreasing sequence which is not bounded above diverges to +infinity.

Proof: Let "\left\{a_{n}\right\}"  be a monotonically non-decreasing sequence.

"\Rightarrow \quad a_{n+1}>a_{n} \forall n \in {N}"

"\left\{a_{n}\right\}" is not bounded above.

"\Rightarrow \quad" For any "M \in {R}, \exists m \in {N}" such that "a_{k}>M \forall k>m"

"\Rightarrow \quad \lim _{n \rightarrow \infty} a_{n}=+\infty"

"\Rightarrow \quad" A monotonically increasing sequence that is not bounded above diverges to positive infinity.